Circular No. 8137

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003em IN ESO 478-G6 Further to IAUC 8016, L. A. G. Monard reports his discovery of an

apparent supernova (R = 17.2) on unfiltered CCD images taken on May 27.158 UT (confirmed at R = 16.7 on May 28.135). The new object is located at α = 2^{h}09^{m}20^{s}.23, δ = −23^{o}24′53′′.0 (equinox 2000.0), which is 32′′ east and 2′′ north of the nucleus of ESO 478-G6. SN 2003em is absent from previous Monard images (Mar. 10.7 and earlier) and a red DSS image.

SUPERNOVA 1988ai R. Evans, Hazelbrook, N.S.W., reports his discovery of an apparent

supernova (mag ∼ 16) on a European Southern Observatory red-survey film taken on 1988 Nov. 7. R. H. McNaught, Siding Spring Observatory, provides the following measurement for SN 1988ai from the discovery film: α = 0^{h}06^{m}21^{s}.74, δ = −41^{o}28′59′′.9 (equinox 2000.0), which is near the northeastern edge of the apparent irregular host galaxy, ESO 293-G34. S. Tritton, Royal Observatory, Edinburgh, reports that the object also appears on a U.K. Schmidt Telescope plate taken on 1989 Feb. 8, but it is not present on other plates of this field (including the J or B Surveys).

η CARINAE

M. F. Corcoran, Universities Space Research Association and Goddard Space Flight Center, reports: “A May 26 observation by the Rossi X-ray Timing Explorer Proportional Counter Array (RXTE PCA) of the peculiar luminous star η Car shows that the 2–10-keV x-ray flux was ≈ 2.7 × 10^{−10} erg s^{−1} cm^{−2}, corresponding to an x-ray luminosity of ∼ 6.5 × 10^{34} erg/s for a distance of 2100 pc. This is the highest x-ray flux yet observed for η Car. Monitoring observations with the RXTE PCA show that the star has undergone four x-ray maxima since Mar. 2003. Three of the x-ray peaks are separated by ∼ 22 days, though the shortest separation is only 12 days. A straight-line fit to three of the four recent peaks, and a linear extrapolation of the rise in brightness since the most recent x-ray minimum on May 9, suggests that the x-ray flux will reach a maximum near 3 × 10^{−10} erg s^{−1} cm^{−2} near May 31. Based on η Car’s 2020-day periodicity (Damineli et al. 2000, Ap.J. 528, L101), this will probably be the maximum brightness of the star during this current cycle, and we expect that the x-ray eclipse ingress (Ishibashi et al. 1999, Ap.J. 524, 983) to begin shortly thereafter. If so, and if the x-ray flux declines at a similar rate to the 1997 eclipse ingress, then the x-ray minimum will occur near 2003 July 24.”

2003 May 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green